# Physical properties of strong 1 ¡ z ¡ 3 Balmer and Paschen lines emitters observed with JWST

### v1

@@ Abstract
Context. The ultraviolet continuum traces young stars while the near-infrared unveils older stellar populations and dust-obscured regions. Balmer emission lines provide insights on gas properties and young stellar objects but are highly affected by dust attenuation. The near-infrared Paschen lines suffer less dust attenuation and can be used to measure star formation rates (SFRs) in star-forming regions obscured by dust clouds.
@@
Methods. We select 13 sources between redshifts 1 and 3 observed with HST, JWST/NIRCam and NIRSpec based on the availability of at least one Balmer and one Paschen line with S/N 5. With a newly-developed version of CIGALE, we fit their hydrogen line equivalent widths (EWs) and photometric data. We assess the impacts of the removal of spectroscopic data by comparing the quality of the fits of the spectro-photometric data to those with photometric data only. We compare the single (BC03) vs binary (BPASS) stellar populations models in the fitting process of spectro-photometric data. We derive the differential attenuation and explore different attenuation recipes by fitting spectro-photometric data with BC03. For each stellar model and for each input dataset (with and without EWs), we quantify the deviation on the SFRs and stellar masses from the ”standard” choice.
Results. The combination of spectro-photometric data provides robust constraints on the physical properties of galaxies, with a significant reduction in the uncertainties compared to using only photometric data. On average, the SFRs are overestimated and the stellar masses are underestimated when EWs are not included as input data. We find a major contribution of the H emission line to the broadband photometric measurements of our sources, and a trend of increasing contribution with specific SFR. Using the BPASS models has a significant impact on the derived SFRs and stellar masses, with SFRs being higher by an average of 0.34 dex and stellar masses being lower by an average of 0.18 dex compared to BC03. We show that a flexible attenuation recipe provides more accurate estimates of the dust attenuation parameters, especially the differential attenuation which agrees with the original value of Charlot & Fall (2000). Finally, we reconstruct the total effective attenuation curves of the most dust-obscured galaxies in our sample.
@@
The ultraviolet (UV) to near-infrared (NIR) spectral energy distribution enables the characterisation of stellar age distributions and the assessment of star formation histories (SFHs) of galaxies (e.g. Salim et al., 2005, 2007). The UV continuum is a tracer of young stars, while the infrared (IR) unveils the older stellar populations and dust-obscured regions within galaxies. Dust grains preferentially absorb shorter-wavelength light, leading to dimming effects that are prominent in the UV and optical regions. In contrast, the NIR is much less affected by dust attenuation making it essential to derive unbiased star formation rates (SFRs) and stellar masses (e.g. Calzetti et al., 2000; Leja et al., 2017). Emission lines, such as those arising from ionised hydrogen, provide crucial insights on gas properties and very young ( 10 Myrs) stellar objects (Kennicutt, 1998; Boselli et al., 2009). A wide spectral coverage of both continuum and emission line features thus enables robust constraints on SFRs, gas metallicity, and ionisation conditions (Shivaei et al., 2020; Reddy et al., 2023).
@@
For the last two decades, astronomers have gathered data from ground-based multi-object NIR spectrographs, such as Keck/MOSFIRE and VLT/KMOS, and space-based facilities like the HST/WFC3 grism, which provided extensive measurements of hydrogen recombination lines and nebular reddening for thousands of galaxies up to a redshift of z 2.6 (e.g. Kashino et al., 2013; Reddy et al., 2015; Shivaei et al., 2020; Rezaee et al., 2021; Battisti et al., 2022). While using these instruments, only a handful of Paschen lines have been robustly detected in z ¿ 2 lensed galaxies (Papovich et al., 2009; Finkelstein et al., 2011). With the advent of the James Webb Space Telescope (JWST, Gardner et al., 2006, 2023) equipped with its NIRSpec instrument, Paschen lines are now becoming accessible for intermediate redshift sources ( is detectable until z 1.7 and until z 3).
@@ 2 Data and sample
The data used in this work are taken from the Cosmic Evolution Early Release Science Survey (CEERS; ERS 1345, PI: S. Finkelstein) in the CANDELS Extended Groth Strip (EGS, Davis et al. (2007); Noeske et al. (2007)) field. The main features of the CEERS program are presented in Arrabal Haro et al. (2023b) and Finkelstein et al. (2022), and will be described in more detail in Finkelstein et al. (in prep.). For each object in this study, we need spectroscopic data to trace the hydrogen line emission of the gas ionised by the very young stars and as much photometric data as we can gather to trace the stellar continuum flux.
@@ 2.1 Spectroscopy
These NIRSpec observations are split into 6 different MSA pointings, each of them observed with the G140M/F100LP, G235M/F170LP, and G395M/F290LP medium resolution () gratings. The total wavelength coverage of the NIRSpec instrument is from 1 to 5.2 m. The MSA is configured to use 3-shutter slitlets, enabling a 3-point nodding pattern, shifting the pointing by a shutter length plus the size of the bar between the shutters in each direction for background subtraction. The total exposure time per grating is 3107s for each MSA pointing.
@@
The first step is to correct the raw images for the detector 1/f noise, dark current, and bias, and ”snowball” contaminations (trails produced by cosmic rays). Count-rate maps are then used to create two-dimensional (2D) spectra for each source.
The second step is to substract the background by using the 3-point nodding pattern, apply the flat-field correction, and calibrate the photometry and wavelength of these spectra. They are then resampled in case of spectral trace distortions. The multiple background subtracted 2D spectra (one for each nod position) are then combined to create the final 2D spectrum of the source.
A post-combination inspection check is done on the reduced 2D spectra in which hot pixels, detector gaps, and reduction artefacts are masked. The noise spectrum is automatically calculated by the JWST pipeline, and then corrected as described in Arrabal Haro et al. (2023a) to take into account the spectral resampling. In the pipeline, a slit loss correction is performed by default in the pathloss step.
The line fluxes and EWs are measured using the beta release of LiMe (Fernández et al., 2023, 2024, submitted). The fluxes are measured assuming a Gaussian distribution and taking into consideration the pixel uncertainty produced by the NIRSpec reduction described in Arrabal Haro et al. (2023a) for the CEERS 0.7 data release. The reported flux uncertainty comes from the standard error definition of the least-squares trust-region method. This uncertainty is propagated to the equivalent width calculation. In this computation, the linear continuum flux and its standard deviation, is derived from two adjacent continua bands (see Fig 1 for an example of line fitting).
These continua regions are manually adjusted for every line, to guarantee a representative linear continuum measurement and to avoid artefacts. We took special care to confirm that the pixel uncertainty was of the same order than the flux standard deviation from the adjacent continua. The LiMe fittings for the 0.7 release spectra used in this work can be found at the interactive repository, 222https://ceers-data.streamlit.app/ using the corresponding NIRSpec MSA IDs. This site will be updated as the JWST pipeline and flux calibration keep improving. Additionally, this manuscript line measurements can be found on the supplementary online material as a FITS file, where each page contains the measurements from one galaxy. The complete measurements description can be found in the online documentation. 333https://lime-stable.readthedocs.io/en/latest/index.html
@@ 2.2 Photometry
Our sample galaxies were also observed with JWST/Near Infrared Camera (NIRCam, Rieke et al., 2023) in December 2022. The 10 pointings and the reduction steps are described in Bagley et al. (2022). We adopt the multi-band SExtractor catalog described in Finkelstein et al. (2022). This catalog includes all available HST/ACS and WFC3 data (v1.9) and all seven JWST/NIRCam bands (v0.51).
@@
Fluxes from the Multiband Imaging Photometer for Spitzer/MIPS, 24m band, the 450 and 850 m James Clerk Maxwell Telescope (JCMT)/SCUBA-2 bandswere retrieved from the Super-deblended Catalog of the EGS field of Le Bail et al. (2023).
@@ 2.3 Sample selection
Galaxies included in our study were selected based on several criteria. The first criterion is to have data from NIRCam and NIRSpec gratings since the spatial coverage of both instruments does not completely overlap (this gives us 126 sources). As we also combine NIRCam data with photometry obtained by other facilities, we need all of them to be in agreement. We checked the overlaps between HST/NIRCam to only keep galaxies for which fluxes measured in the same aperture of nearby bands (HST/F160W and NIRCam/F150W) were within (122 objects remain at this stage).We present the photometric data used in the study in Table 1.
Balmer and Paschen lines are necessary as they probe stars located in vastly different environments so we require our objects to have at least one Balmer and one Paschen line with an equivalent width S/N (other Balmer and Paschen lines with a S/N are also added, see Table 2). Getting the lines to compute the ratio of a Paschen line to a Balmer line (Pa/H hereafter) described in Section 3.1 restricts our sample to the redshift range 1 ¡ z ¡ 3. All CEERS spectra were inspected to identify and secure a spectroscopic redshift between 1 and 3 which amounts to 40 unique sources. The requirement of a S/N for at least one Balmer line and one Paschen line further restricts our sample down to 17 objects.
Finally, we want to avoid galaxies with an active galactic nucleus (AGN) as their presence would require much more complex models beyond the scope of this paper. To check if our sources could host AGNs, we compute the ratio of the [NII]-6585 over H if the [NII]-6585 line is detected (a non-detection of the [NII]-6585 line is indicated by a cross in Table 2). [NII]-6585/H is a linear function of the nebular metallicity and it presents a saturation point so a further increase in the [NII]-6585/H value is only due to AGN contribution (Stasińska et al., 2006). Following the authors’ prescription, we only keep galaxies with log([NII]-6585/H) ¡ -0.2 (4 objects were dropped at this stage). Some sources (IDs 8515, 8588 and 8710) are considered ”hidden AGNs” by Calabrò et al. (2023) following their diagnostics based on IR emission lines. These AGNs diagnostics are BPT-like diagrams (Baldwin et al., 1981) of ratios of [CI]-9850, [PII] 1.188m or [FeII] 1.257m to Paschen lines (either Pa or Pa) versus a ratio of [SIII]-9530 to Pa or Pa. As we do not detect [PII] or [CI] in these 3 galaxies and two of them have a [FeII] detection with a S/N ¡ 3 ([FeII] is not detected for 8588), we decided to still include them as their exact nature as AGNs still needs to be confirmed. Using all of these criteria, we end up with 13 galaxies in our sample as requiring an observation with NIRCam and the NIRSpec gratings drastically reduces the number of available sources.
@@ 3.1 Theoretical line ratios
Within our sample, we are able to get multiple Balmer and Paschen lines. In this section, we will compute theoretical ratios of lines, i.e. ratios expected for a case B recombination and a temperature of 10000 (Osterbrock, 1989), as well as model the reddening of the stellar emission following a CF00 attenuation recipe. The key feature of this model is the computation of two complementary recipes, one for the birth clouds (BCs) and one for the interstellar medium (ISM) In this way, we are able to properly account for an age-dependent attenuation where not only the total amount of dust attenuation changes as a function of the stellar age but also the way the stellar light is obscured, as originally introduced by CF00.
In this model, young stars up to an age of 10 Myrs are surrounded by BCs which progressively dissipate over time. Dust attenuation works differentially, with an efficient attenuation in the BCs combined with a lower attenuation in the surrounding ISM. The double component attenuation CF00 recipe is critical as it means that the emission lines and UV light emitted by young stars are attenuated by both the BCs and the ISM. However, once stars have migrated outside the BCs, the emitting radiation is subject only to the diffuse ISM attenuation. Throughout this study, we will refer to the results or models of dust attenuation as ”recipes” and not ”curves”. Indeed, the latter would suggest that a unique, universal result exists while the real effective result depends on the SFH and attenuation parameters of each individual galaxy.
@@
where is the slope of the attenuation curve for the ISM and the slope of the attenuation curve for the birth clouds. Both power-law exponents are fixed to -0.7 in the original recipe of CF00. However, these parameters have been shown to vary among galaxies (e.g. Buat et al., 2012; Chevallard et al., 2013; Kriek & Conroy, 2013; Battisti et al., 2017; Lo Faro et al., 2017; Salim et al., 2018; Trayford et al., 2020; Pantoni et al., 2021; Boquien et al., 2022) so we decided to keep these exponents as free parameters. The attenuation for the birth clouds, , is computed using the parameter:
@@
We use the Balmer decrement (H/H) as it is available for all our sources with fluxes measured with a high S/N. For the Pa/H lines ratio, we have many different combinations available thanks to the large wavelength coverage of the three gratings. We check which combination gives us the best chance at discriminating between two different exponents for a power law which follows Eq 1 of the CF00 recipe by plotting the Pa/H ratios at our disposal in Fig 2. The two ratios with the widest gap are Pa/H and Pa/H, they are however, only available for a few sources in our sample (see Table 2). Because of this issue, we settle for the two next best ratios: Pa/H and Pa/H (in cases where both ratios are available, both will be included in the analysis). Finally we note that the Pa/H ratio needs very high levels of attenuation to start being discriminant between both exponents, so we will not be using this ratio in this section.

### v3

# Physical properties of strong 1 ¡ z ¡ 3 Balmer and Paschen line emitters observed with JWST

@@ Abstract
Context. The ultraviolet continuum traces young stars while the near-infrared unveils older stellar populations and dust-obscured regions. Balmer emission lines provide insights into gas properties and young stellar objects but are highly affected by dust attenuation. The near-infrared Paschen lines suffer less dust attenuation and can be used to measure star formation rates (SFRs) in star-forming regions obscured by dust clouds.
@@
Methods. We select 13 sources between redshifts 1 and 3 observed with HST, JWST/NIRCam and NIRSpec based on the availability of at least one Balmer and one Paschen line with S/N 5. With a newly developed version of CIGALE, we fit their hydrogen line equivalent widths (EWs) and photometric data. We assess the impacts of the removal of spectroscopic data by comparing the quality of the fits of the spectro-photometric data to those with photometric data only. We compare the single (BC03) vs binary (BPASS) stellar population models in the fitting process of spectro-photometric data. We derive the differential attenuation and explore different attenuation recipes by fitting spectro-photometric data with BC03. For each stellar model and for each input dataset (with and without EWs), we quantify the deviation on the SFRs and stellar masses from the ”standard” choice.
Results. The combination of spectro-photometric data provides robust constraints on the physical properties of galaxies, with a significant reduction in the uncertainties compared to using only photometric data. On average, the SFRs are overestimated and the stellar masses are underestimated when EWs are not included as input data. We find a major contribution of the H emission line to the broadband photometric measurements of our sources, and a trend of increasing contribution with specific SFR. Using the BPASS models has a significant impact on the derived SFRs and stellar masses, with SFRs being higher by an average of 0.13 dex and stellar masses being lower by an average of 0.18 dex compared to BC03. We show that a flexible attenuation recipe provides more accurate estimates of the dust attenuation parameters, especially the differential attenuation. Finally, we reconstruct the total effective attenuation curves of the most dust-obscured galaxies in our sample.
@@
The ultraviolet (UV) to near-infrared (NIR) spectral energy distribution enables the characterisation of stellar age distributions and the assessment of star formation histories (SFHs) of galaxies (e.g. Salim et al., 2005, 2007). The UV continuum is a tracer of young stars, while the infrared (IR) unveils the older stellar populations and dust-obscured regions within galaxies. Dust grains preferentially absorb shorter-wavelength light, leading to dimming effects that are prominent in the UV and optical regions. In contrast, the NIR is much less affected by dust attenuation making it essential to derive unbiased star formation rates (SFRs) and stellar masses (e.g. Calzetti et al., 2000; Leja et al., 2017). Emission lines, such as those arising from ionised hydrogen, provide crucial insights into gas properties and very young ( 10 Myrs) stellar objects (Kennicutt, 1998; Boselli et al., 2009). A wide spectral coverage of both continuum and emission line features thus enables robust constraints on SFRs, gas metallicity, and ionisation conditions (Shivaei et al., 2020; Reddy et al., 2023).
@@
For the last two decades, astronomers gathered data from ground-based multi-object NIR spectrographs, such as Keck/MOSFIRE and VLT/KMOS, and space-based facilities like the HST/WFC3 grism, which provided extensive measurements of hydrogen recombination lines and nebular reddening for thousands of galaxies up to a redshift of z 2.6 (e.g. Kashino et al., 2013; Reddy et al., 2015; Shivaei et al., 2020; Rezaee et al., 2021; Battisti et al., 2022). While using these instruments, only a handful of Paschen lines have been robustly detected in z ¿ 2 lensed galaxies (Papovich et al., 2009; Finkelstein et al., 2011). With the advent of the James Webb Space Telescope (JWST, Gardner et al., 2006, 2023) equipped with its NIRSpec instrument, Paschen lines are now becoming accessible for intermediate redshift sources ( is detectable until z 1.7 and until z 3).
@@ 2 Data and sample
The data for this work are taken from the Cosmic Evolution Early Release Science Survey (CEERS; ERS 1345, PI: S. Finkelstein) in the CANDELS survey (Grogin et al., 2011; Koekemoer et al., 2011) of the Extended Groth (EGS, Davis et al. (2007); Noeske et al. (2007)) field. The main features of the CEERS program are presented in Arrabal Haro et al. (2023b) and Finkelstein et al. (2023), and will be described in more detail in Finkelstein et al. (in prep.). For each object in this study, we need spectroscopic data to trace the hydrogen line emission of the gas ionised by the very young stars and as much photometric data as we can gather to trace the stellar continuum flux.
@@ 2.1 Spectroscopy
These NIRSpec observations are split into six different MSA pointings, each of them observed with the G140M/F100LP, G235M/F170LP and G395M/F290LP medium resolution (R = 1000) gratings. The total wavelength coverage of the NIRSpec instrument is from 1 to 5.2 m. The MSA is configured to use three shutter slitlets thus enabling a three point nodding pattern, shifting the pointing by a shutter length plus the size of the bar between the shutters in each direction for background subtraction. The total exposure time per grating is 3107s for each MSA pointing.
@@
The first step is to correct the raw images for the detector 1/f noise, dark current, bias and ”snowball” contaminations (trails produced by cosmic rays). Count-rate maps are then used to create two-dimensional (2D) spectra for each source.
The second step is to subtract the background by using the three point nodding pattern, apply the flat-field correction, and calibrate the photometry and wavelength of these spectra. They are then resampled in case of spectral trace distortions. The multiple background subtracted 2D spectra (one for each nod position) are then combined to create the final 2D spectrum of the source.
A post-combination inspection check is done on the reduced 2D spectra in which hot pixels, detector gaps and reduction artefacts are masked. The noise spectrum is automatically calculated by the JWST pipeline and then corrected as described in Arrabal Haro et al. (2023a) to take into account the spectral resampling. In the pipeline a slit loss correction is performed by default in the pathloss step.
The line fluxes and EWs are measured using the beta release of LiMe (Fernández et al., 2023, 2024, accepted for publication in A&A). The fluxes are measured assuming a Gaussian distribution and taking into consideration the pixel uncertainty produced by the NIRSpec reduction described in Arrabal Haro et al. (2023a) for the CEERS 0.7 data release. The reported flux uncertainty comes from the standard error definition of the least-squares trust-region method. This uncertainty is propagated to the equivalent width calculation. In this computation the linear continuum flux and its standard deviation are derived from two adjacent continua bands (see Fig 1 for an example of line fitting).
These continua regions are manually adjusted for every line to guarantee a representative linear continuum measurement and to avoid artefacts. We take special care to confirm that the pixel uncertainty was of the same order than the flux standard deviation from the adjacent continua. The LiMe fittings for the 0.7 release spectra used in this work can be found at the interactive repository 222https://ceers-data.streamlit.app/ using the corresponding NIRSpec MSA IDs. This site will be updated as the JWST pipeline and flux calibration keep improving. Additionally, this manuscript line measurements can be found on the supplementary online material as a FITS file, where each page contains the measurements from one galaxy. The complete measurements description can be found in the online documentation. 333https://lime-stable.readthedocs.io/en/latest/index.html
@@ 2.2 Photometry
Our sample galaxies were also observed with JWST/Near Infrared Camera (NIRCam, Rieke et al., 2023) in December 2022. The ten pointings and the reduction steps are described in Bagley et al. (2023). We adopt the multi-band SExtractor catalog described in Finkelstein et al. (2023). This catalog includes all available HST/ACS and WFC3 data (v1.9) and all seven JWST/NIRCam bands (v0.51).
@@
Fluxes from the Multiband Imaging Photometer for Spitzer/MIPS, 24m band, the 450 and 850 m James Clerk Maxwell Telescope (JCMT)/SCUBA-2 bandsare retrieved from the Super-deblended Catalog of the EGS field of (Le Bail et al., 2023, accepted for publication in A&A ).
@@ 2.3 Sample selection
Galaxies included in our study are selected based on several criteria. The first criterion is to have data from NIRCam and NIRSpec gratings since the spatial coverage of both instruments does not completely overlap (this gives us 126 sources). As we also combine NIRCam data with photometry obtained by other facilities, we need all of them to be in agreement. The overlaps between HST/NIRCam are checked to only keep galaxies for which fluxes measured in the same aperture of nearby bands (HST/F160W and NIRCam/F150W) were within (122 objects remain at this stage). We present the photometric data used in the study in Table 4.
Balmer and Paschen lines are necessary as they probe stars located in vastly different environments so we require our objects to have at least one Balmer and one Paschen line with an equivalent width S/N (other Balmer and Paschen lines with a S/N are also added, see Table 5). Getting the lines to compute the ratio of a Paschen line to a Balmer line (Pa/H hereafter) described in Section 3.1 restricts our sample to the redshift range 1 ¡ z ¡ 3. All CEERS spectra are inspected to identify and secure a spectroscopic redshift between 1 and 3 which amounts to 40 unique sources. The requirement of a S/N for at least one Balmer line and one Paschen line further restricts our sample down to 17 objects.
Finally, we want to avoid galaxies with an active galactic nucleus (AGN) as their presence would require much more complex models beyond the scope of this paper. To check if our sources could host AGNs, we compute the ratio of the [NII]-6585 over H if the [NII]-6585 line is detected (a non-detection of the [NII]-6585 line is indicated by a cross in Table 5). [NII]-6585/H is a linear function of the nebular metallicity and it presents a saturation point so a further increase in the [NII]-6585/H value is only due to AGN contribution (Stasińska et al., 2006). Following the authors’ prescription, we only keep galaxies with log([NII]-6585/H) ¡ -0.2 (four objects were dropped at this stage). Some sources (IDs 8515, 8588 and 8710) are considered ”hidden AGNs” by Calabrò et al. (2023) following their diagnostics based on IR emission lines. These AGNs diagnostics are BPT-like diagrams (Baldwin et al., 1981) of ratios of [CI]-9850, [PII] 1.188m or [FeII] 1.257m to Paschen lines (either Pa or Pa) versus a ratio of [SIII]-9530 to Pa or Pa. As we do not detect [PII] or [CI] in these three galaxies and two of them have a [FeII] detection with a S/N ¡ 3 ([FeII] is not detected for 8588), we decide to still include them as their exact nature as AGNs still needs to be confirmed. Using all of these criteria, we end up with 13 galaxies in our sample as requiring an observation with NIRCam and the NIRSpec gratings drastically reduces the number of available sources.
@@ 3.1 Theoretical line ratios
Within our sample, we are able to get multiple Balmer and Paschen lines. In this section, we compute theoretical ratios of lines, i.e. ratios expected for a case B recombination and a temperature of 10 000 (Osterbrock, 1989), as well as model the reddening of the stellar emission following a CF00 attenuation recipe. The key feature of this model is the computation of two complementary recipes, one for the birth clouds (BCs) and one for the interstellar medium (ISM) In this way, we are able to properly account for an age-dependent attenuation where not only the total amount of dust attenuation changes as a function of the stellar age but also the way the stellar light is obscured, as originally introduced by CF00.
In this model, young stars up to an age of 10 Myrs are surrounded by BCs which progressively dissipate over time. Dust attenuation works differentially, with an efficient attenuation in the BCs combined with a lower attenuation in the surrounding ISM. The double component attenuation CF00 recipe is critical as it means that the emission lines and UV light emitted by young stars are attenuated by both the BCs and the ISM. However, once stars have migrated outside the BCs, the emitting radiation is subject only to the diffuse ISM attenuation. Throughout this study, we refer to the results or models of dust attenuation as ”recipes” and not ”curves”. Indeed, the latter would suggest that a unique or universal result exists while the real effective result depends on the SFH and attenuation parameters of each individual galaxy.
@@
where is the slope of the attenuation curve for the ISM and the slope of the attenuation curve for the birth clouds. Both power-law exponents are fixed to -0.7 in the original recipe of CF00. However, these parameters have been shown to vary among galaxies (e.g. Buat et al., 2012; Chevallard et al., 2013; Kriek & Conroy, 2013; Battisti et al., 2017; Lo Faro et al., 2017; Salim et al., 2018; Trayford et al., 2020; Pantoni et al., 2021; Boquien et al., 2022) so we decide to keep these exponents as free parameters. The attenuation for the birth clouds, , is computed using the parameter:
@@
We use the Balmer decrement (H/H) as it is available for all our sources with fluxes measured with a high S/N. For the Pa/H lines ratio, we have many different combinations available thanks to the large wavelength coverage of the three gratings. We check which combination gives us the best chance at discriminating between two different exponents for a power law which follows Eq 1 of the CF00 recipe by plotting the Pa/H ratios at our disposal in Fig 2. The two ratios with the widest gap are Pa/H and Pa/H, they are however, only available for a few sources in our sample (see Table 5). Because of this issue, we settle for the two next best ratios: Pa/H and Pa/H (in cases where both ratios are available, both will be included in the analysis). Finally we note that the Pa/H ratio needs very high levels of attenuation to start being discriminant between both exponents, so we are not using this ratio in this section.